# Evidence of Energy Injection in the Short and Distant GRB 250221A

### v4

@@ Appendix B Central Engine
In order to produce a GRB we need a central engine, whether the event is coming from a compact-objects merger, i.e., a low mass BH-NS, or a NS-NS (Lattimer & Schramm, 1974; Metzger et al., 2011)–, or from a Collapsar as was explained in Section 1 (Woosley, 1993). So far, neither our restricted observations nor our analytical models can tell which one is responsible for GRB 250221A.
@@
    $\nu\bar{\nu}$ annihilation from accretion onto a NS or BH (see, e.g. MacFadyen & Woosley, 1999b; Zalamea & Beloborodov, 2011; Globus & Levinson, 2014): When the mass-accretion rate close to the compact object (CO: BH or NS) is larger than a few hundredths of solar masses per second and the spin (of the CO) is large; or accretion rates around a few tenths of a solar mass per second and the spin is low; conditions are optimal for nuclear burning in the inner disk and, thus, production of large amounts of $\nu\bar{\nu}$. Zalamea & Beloborodov (2011) have shown that powers $\dot{E}_{\nu\bar{\nu}}\gtrsim 10^{52}$ erg s-1 can be reached for accretion rates $\gtrsim 1$ M s-1, and cannot exceed $\sim 5\,\times 10^{52}$ erg s-1.
@@
    A Kerr BH with Blandford-Znajek engine (Blandford & Znajek, 1977; Brown et al., 2000): A Penrose process (Penrose & Floyd, 1971) where a Kerr BH, with an attached magnetic field finds itself surrounded by a plasma-built accretion disk, works as an electric generator and the average of the Poynting vector points in the direction of the rotational axis, delivering thus, a fraction of the Kerr’s BH rotational kinetic energy into jets. Typical Blandford-Znajek (BZ) luminosities ($L_{BZ}$) are of order $L_{BZ}\simeq 10^{52}a_{\star}^{2}\frac{\Omega(\omega_{BH}-\Omega)}{\omega_{BH}^{2}}\left(\frac{M_{BH}}{5M_{\odot}}\right)^{2}\left<\frac{B}{10^{15}{\rm G}}\right>^{2}$ erg s-1, where $a_{\star}$ is the dimensionless Kerr (or spin) parameter of the BH, $\omega_{BH}$ is the angular velocity of the BH, and $\Omega$ is the angular velocity of the accretion disk.
@@
    Schwarschild BH or Kerr BH with BP engine (Blandford & Payne, 1982): Again, an electric-generator like process, but not a Penrose process. This is magneto-rotational process that extracts rotational energy from the accretion disk.

### v5

@@ Appendix B Central Engine
In order to produce a GRB we need a central engine, whether the event is coming from a compact-objects merger, i.e., a low mass BH-NS, or a NS-NS (Lattimer and Schramm, 1974; Metzger et al., 2011)–, or from a Collapsar as was explained in Section 1 (Woosley, 1993). So far, neither our restricted observations nor our analytical models can tell which one is responsible for GRB 250221A.
@@
    $\nu\bar{\nu}$ annihilation from accretion onto a NS or BH (see, e.g. MacFadyen and Woosley, 1999a; Zalamea and Beloborodov, 2011; Globus and Levinson, 2014): When the mass-accretion rate close to the compact object (CO: BH or NS) is larger than a few hundredths of solar masses per second and the spin (of the CO) is large; or accretion rates around a few tenths of a solar mass per second and the spin is low; conditions are optimal for nuclear burning in the inner disk and, thus, production of large amounts of $\nu\bar{\nu}$. Zalamea and Beloborodov (2011) have shown that powers $\dot{E}_{\nu\bar{\nu}}\gtrsim 10^{52}$ erg s-1 can be reached for accretion rates $\gtrsim 1$ M s-1, and cannot exceed $\sim 5\,\times 10^{52}$ erg s-1.
@@
    A Kerr BH with Blandford-Znajek engine (Blandford and Znajek, 1977; Brown et al., 2000): A Penrose process (Penrose and Floyd, 1971) where a Kerr BH, with an attached magnetic field finds itself surrounded by a plasma-built accretion disk, works as an electric generator and the average of the Poynting vector points in the direction of the rotational axis, delivering thus, a fraction of the Kerr’s BH rotational kinetic energy into jets. Typical Blandford-Znajek (BZ) luminosities ($L_{BZ}$) are of order $L_{BZ}\simeq 10^{52}a_{\star}^{2}\frac{\Omega(\omega_{BH}-\Omega)}{\omega_{BH}^{2}}\left(\frac{M_{BH}}{5M_{\odot}}\right)^{2}\left<\frac{B}{10^{15}{\rm G}}\right>^{2}$ erg s-1, where $a_{\star}$ is the dimensionless Kerr (or spin) parameter of the BH, $\omega_{BH}$ is the angular velocity of the BH, and $\Omega$ is the angular velocity of the accretion disk.
@@
    Schwarschild BH or Kerr BH with BP engine (Blandford and Payne, 1982): Again, an electric-generator like process, but not a Penrose process. This is magneto-rotational process that extracts rotational energy from the accretion disk.